Explain this example of the following water quality serial dilution. Note the dilution factors and how you would calculate #;s from the incubated plates. A serial dilution yielded TBTC in the 1:1 plate and the 1:10 plate, 10 cfu’s in the 1:100 plate and 1 cfu’s in the 1:1000 plate. What is the concentration (CFU)/ml of the raw sample? Is water acceptable for portable or Recreational use? Explain. When do you use your filtration rather than serial dilution?
Answer #1

Serial dilution is done when we assume that first dilution does have high load i.e. TNTC, which can not be counted. To count the CFU/ml, we go further dilution until we get CFU which can be count easily. Then, CFU/ml can be count = CFU * serial dilution factor.

In calculation via conventional formula,CFU/ml in 2nd dilution are greater than the 1st dilution)

For example

D1; 689 CFU/0.1 ml

CFU/ml= 689*10/0.1=6890

D2: 478 CFU/0.1 ml

CFU/ml= 478*100/0.1= 478000

When we do a serial dilution, it means that we are reducing the concentration of the cells/microbes so that it can be easily counted. At a higher dilution (depending on the microbes concentration in the original sample) we can see few and separated colonies so that we count it, multiply it by the dilution factor like, but it should also be divided to the amount of volume we poured onto the plate in order to get CFU/ml. This doesn't mean that the CFU/ml should decrease.

CFU per ml does not depend on Dilution,

think first dilution give 80 colony then 80*10 = 800 CFU/ml

then second dilution gives 8 colonies then 8*100 = 800 CFU/ml

when we are testing for the probiotic we take 10-8 ,10-9 , 10-10 of dilution then our results followed,

>>10-8

colony gives = 396

then 396*108=3.96*1010 CFU/g

>>>10-9

Colony gives = 40

then 40*109=4.0*1010 CFU/g

>>>10-10

Colonies = 3 to 5 then mean is 4

then 4*1010 CFU/g

from the above, you can think dilution does not affect the CFU/ml,

we have to dilute our sample to count the colonies easily.

we count those plates who has count b/w 10-300. above 300 per plate colonies discarded and written as TNPC= To Numeric Plate Count.